OpenStudy (anonymous):

Write an equation of the line, in point-slope form, that passes through the two given points. points: (–2, 15), (9, –18)

OpenStudy (radar):

This may help. m = slope. \[x _{1},y _{1}= P1\]\[x _{2},y _{2}= P _{2}\]\[m=(y _{2}-y _{1})/(x _{2}-x _{1})\]

OpenStudy (anonymous):

so the answer is .. y – 15 = –3(x + 2) ?

OpenStudy (radar):

First step find slope m: Let me see if you are on the right track.

OpenStudy (anonymous):

slope = -3

OpenStudy (radar):

m=(15 - (-18)) /(-2-9) m=33/-11=-3 That is what you have. Is there a requirement to what form the equation must be ? Slope Intercept or Standard? You have basically solved the problem.

OpenStudy (anonymous):

point slope form

OpenStudy (radar):

Slope intercept form would be: y=-3x - 6 + 15 or y=-3x + 9

OpenStudy (radar):

I think you have it in the point slope form.

OpenStudy (radar):

Yes your answer was in the correct point slope form according to my book.

OpenStudy (anonymous):

thanks

OpenStudy (radar):

Point slope form is: y-y1 = m(x-x1) Great work and good luck with your studies.